# ConceptCaps: a Distilled Concept Dataset for Interpretability in Music Models

### v2

@@ Abstract
Concept-based interpretability methods like TCAV require clean, well-separated positive and negative examples for each concept. Existing music datasets lack this structure: tags are sparse, noisy, or ill-defined. We introduce ConceptCaps, a dataset of 21k music-caption-audio triplets with explicit labels from a 200-attribute taxonomy. Our pipeline separates semantic modeling from text generation: a VAE learns plausible attribute co-occurrence patterns, a fine-tuned LLM converts attribute lists into professional descriptions, and MusicGen synthesizes corresponding audio. This separation improves coherence and controllability over end-to-end approaches. We validate the dataset through audio-text alignment (CLAP), linguistic quality metrics (BERTScore, MAUVE), and TCAV analysis confirming that concept probes recover musically meaningful patterns. Dataset and code are available online.
@@ 3.6 Implementation
Source code, datasets and pre-trained models are publicly available online for reproducibility purposes [^1] https://github.com/BrunoSienkiewicz/ConceptCaps. We are using $\beta$-VAE Higgins et al. (2017) variant of variational auto-encoder with additional $\beta$ parameter, which controls KL divergence loss. From our experiments we concluded that $\beta=0.25$ yielded the best overall results. It suggests that minimizing reconstruction loss is more important in our case than mapping latent space to normal distribution. Higher $\beta$ values resulted in model collapse suggested by VAE predicting only the few most popular attribute combinations. VAE is implemented with a standard fully-connected architecture: encoder with DD-1024 hidden units, where $D$ is denotes our aspect taxonomy dimension, mapping to a 256-dimensional latent space, and a symmetric decoder. We use Adam optimizer Kingma (2014) with learning rate $6\times 10^{-5}$ and train for 150 epochs on the 1,890 distilled samples.
@@
Stage 1: VAE Jaccard $\uparrow$ 0.716 Intersection-over-union of reconstructed vs. ground-truth attribute vectors.
@@
Sound-VECaps Yuan and others (2024) 1,660,000 527 classes Yes $\sim$40 4,611

### v3

@@ Abstract
Concept-based interpretability methods like TCAV require clean, well-separated positive and negative examples for each concept. Existing music datasets lack this structure: tags are sparse, noisy, or ill-defined. We introduce ConceptCaps, a dataset of 21k music-caption-tags triplets with explicit labels from a 200-attribute taxonomy. Our pipeline separates semantic modeling from text generation: a VAE learns plausible attribute co-occurrence patterns, a fine-tuned LLM converts attribute lists into professional descriptions, and MusicGen synthesizes corresponding audio. This separation improves coherence and controllability over end-to-end approaches. We validate the dataset through audio-text alignment (CLAP), linguistic quality metrics (BERTScore, MAUVE), and TCAV analysis confirming that concept probes recover musically meaningful patterns. Dataset and code are available online.
@@ 3.6 Implementation
Source code, datasets and pre-trained models are publicly available online for reproducibility purposes [^1] https://github.com/BrunoSienkiewicz/ConceptCaps. We are using $\beta$-VAE Higgins et al. (2017) variant of variational auto-encoder with additional $\beta$ parameter, which controls KL divergence loss. From our experiments we concluded that $\beta=0.25$ yielded the best overall results. It suggests that minimizing reconstruction loss is more important in our case than mapping latent space to normal distribution. Higher $\beta$ values resulted in model collapse suggested by VAE predicting only the few most popular attribute combinations. VAE is implemented with a standard fully-connected architecture: encoder with DD-512 hidden units, where $D$ is denotes our aspect taxonomy dimension, mapping to a 128-dimensional latent space, and a symmetric decoder. We use Adam optimizer Kingma (2014) with learning rate $3\times 10^{-4}$ and train for 110 epochs on the 1,890 distilled samples.
@@
Stage 1: VAE BCE $\downarrow$ 105.21 Binary Cross-Entropy loss of the reconstructed vectors vs. ground-truth attribute vectors.
Jaccard $\uparrow$ 0.716 Intersection-over-union of reconstructed vs. ground-truth attribute vectors.
@@
Sound-VECaps Yuan and others (2024) 1,660,000 527 classes Yes $\sim$30 4,611